# LibMOON: A Gradient-based MultiObjective OptimizatioN Library in PyTorch

### v2

@@ 3.2 MOO solvers
To find Pareto solutions, a straightforward way is to use some aggregation functions $g_{\bm{\lambda}}^{\mathrm{agg}}():\mathbb{R}^{m}\to\mathbb{R}$ to convert a multi-objective problem (MOP) into a single-objective optimization problem (SOP). Aggregation function based methods aim to optimize the following objective function:
$\min_{\bm{\theta}}g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}})),$ (2)
where ${\bm{L}}({\bm{\theta}})$ represents an objective vector and $g_{\bm{\lambda}}^{\mathrm{agg}}()$ is an aggregation function determined by a preference vector ${\bm{\lambda}}$. If $g_{\bm{\lambda}}^{\mathrm{agg}}()$ is decreasing with respect to ${\bm{L}}({\bm{\theta}})$ (i.e., $g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}^{(a)}))<g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}^{(b)}))$ when ${\bm{L}}_{i}({\bm{\theta}}^{(a)})\leq{\bm{L}}_{i}({\bm{\theta}}^{(b)})$ for all $i$ with at least one strict inequality), the optimal solution of Equation 2 corresponds to a Pareto optimal solution of the original MOP (Equation 1[42].
The linear scalarization (LS) function, $g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}))=\sum_{i=1}^{m}\lambda_{i}L_{i}({\bm{\theta}})$, is the most commonly used aggregation function. Other notable functions include the (Smooth) Tchebycheff (Tche) [25, 42], modified Tchebycheff (mTche) [43], Penalty-Based Intersection (PBI) [33], and COSMOS [37]. For detailed formulations, see Appendix A. Aggregation-based methods can be efficiently optimized in LibMOON using single objective backpropagation. Aggregation-based method update the parameter ${\bm{\theta}}$ iteratively as follows:
${\bm{\theta}}\xleftarrow{}{\bm{\theta}}-\eta{\bm{d}}={\bm{\theta}}-\eta\frac{\partial g_{\bm{\lambda}}^{\mathrm{agg}}({\bm{L}}({\bm{\theta}}))}{\partial{\bm{\theta}}},$ (3)
where ${\bm{d}}$ is called an updating direction.
Generalized aggregation function.
In addition to optimizing the single-objective aggregation function, a number of gradient manipulation methods adjust gradients and compute update directions to ensure solutions satisfy specific properties. These methods can be summarized as follows: the update direction ${\bm{d}}$ is a dynamically weighted sum of gradient vectors [18], expressed as ${\bm{d}}=\sum_{i=1}^{m}\tilde{\alpha}_{i}\cdot\nabla L_{i}({\bm{\theta}})$. Equivalently, updating the direction involves optimizing a generalized aggregation function.
$\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))=\sum_{i=1}^{m}\tilde{\alpha}_{i}\cdot L_{i}({\bm{\theta}}).$ (4)
Multiple gradient manipulation methods.
As listed in Table 3, notable methods using gradient manipulations include EPO [14] for finding “exact Pareto solutions” (intersection points of Pareto front and preference vectors), HVGrad [39] for maximizing hypervolume, MGDA-UB [13] for finding directions beneficial to all objectives, MOO-SVGD [19] for diverse solutions, PMGDA [40] for Pareto solutions satisfying specific user requirement, and PMTL [18] for sector-constrained Pareto solutions.
In LibMOON , multiple gradient manipulation methods are implemented by first calculating dynamic weight factors $\tilde{\alpha}_{i}$ and then performing backpropagation on the generalized function $\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))$ to update the parameters of neural network. Those weight factors $\tilde{\alpha}_{i}$ are typically computed by solving a linear programming (LP) problem (e.g., [14][Eq. 24], [40][Eq. 16]), a quadratic programming (QP) problem (e.g., [18][Eq. 14], [13][Eq. 3]), or through other methods like hypervolume gradient estimation [44].
Some methods use user-specific preference vectors as input, making the aggregation function dependent on preference ${\bm{\lambda}}$, while others do not (preference-free optimizers). A summary of whether an optimizer is preference-based or preference-free is provided in Table 3.
@@
LibMOON not only supports first-order optimization when the gradients of those objective functions are known, but also supports zero-order optimization methods with estimated gradients ^$\hat{\nabla}L_{i}({\bm{\theta}})$ using methods such evolutionary strategy (ES) [45].
@@ 3.3 Pareto set learning solvers
Method Property Paper Term ${\bm{A}}$ Term ${\bm{B}}$
EPO-based PSL [16] Exact solutions paper Specific solvers BP
PMGDA-based PSL [40] Specific solutions paper Specific solvers BP
Aggregation-based PSL [16] Minimal aggregation values paper BP BP
Evolutionary PSL [46] Mitigate local minima by ES paper BP ES
LoRA PSL [47, 24, 48] Light model structure paper BP BP
  • BP: backward propagation. ES: evolutionary strategy. Specific solvers include EPO, PMGDA, etc.
@@
In addition to the existing methods for finite Pareto solutions, LibMOON supports Pareto Set Learning (PSL), which trains a model with parameter ${\bm{\phi}}$ to generate an infinite number of approximate Pareto solutions. We denote the model output as ${\bm{\theta}}$ and represent the model as ${\bm{\theta}}_{\bm{\phi}}(\cdot):{\bm{\Delta}}_{m}\to\mathbb{R}^{n}$, where ${\bm{\Delta}}_{m}$ is the $m$-dim preference simplex and $\mathbb{R}^{n}$ is the entire decision space of ${\bm{\theta}}$.
PSL can be deemed as learning a model that find solutions under infinitely many preference well. The loss function of PSL is formulated as follows:
$\min_{\bm{\phi}}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}({\bm{\Delta}}_{m})}\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))),\qquad\theta_{\phi}(\cdot):{\bm{\Delta}}_{m}\xrightarrow{}\mathbb{R}^{n}.$ (5)
The function $\tilde{g}_{\bm{\lambda}}(\cdot):\mathbb{R}^{m}\to\mathbb{R}$ is a generalized aggregation function (as introduced in Equation 4), which converts an $m$-dim objective vector into a scalar based on specific preferences. Since Equation 5 involves preference expectation, all preference-based finite Pareto solvers (Table 3) can serve as basic Pareto solution solvers, while preference-free solvers are unsuitable for PSL.
Figure 2: Model architecture of Pareto set models.
For PSL, we explain the function ${\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))$ with more details as follows: In synthetic problems, the decision variable is directly the model output ${\bm{\theta}}_{\bm{\phi}}$, and ${\bm{\theta}}_{\bm{\phi}}$ is evaluated by the loss function. In multi-task learning (MTL) problems, ${\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}})$ represents the parameters of a target network, and ${\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))$ is evaluated using the dataset $({\bm{x}}_{i},{\bm{y}}_{i})$ with this target network. See Figure 2 for a visualization of these architectures.
The gradient of $\ell_{\mathrm{psl}}$ can be estimated by the chain rule:
@@
$\underbrace{\frac{\partial\ell_{\mathrm{psl}}}{\partial{\bm{\phi}}}}_{1\times D}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}({\bm{\Delta}}_{m})}\underbrace{\frac{\partial\tilde{g}_{\bm{\lambda}}}{\partial{\bm{L}}}}_{{\bm{A}}:(1\times m)}\cdot\underbrace{\frac{\partial{\bm{L}}}{\partial{\bm{\theta}}}}_{{\bm{B}}:(m\times n)}\underbrace{\frac{\partial{\bm{\theta}}}{\partial{\bm{\phi}}}}_{{\bm{C}}:(n\times D)}.$ (6)
@@
The expectation is empirically estimated using a batch of $K$ preferences. The gradient vector ${\bm{A}}=(\tilde{\alpha}_{1},\ldots,\tilde{\alpha}_{m})$ is computed by the finite Pareto solution solver. The gradient matrix ${\bm{B}}$ can be calculated either via backpropagation (when gradients are easily obtained) or using a zero-order method. In Equation 6, ${\bm{C}}$ is always estimated through backpropagation, since ${\bm{\theta}}$ is a continuous vector function of ${\bm{\phi}}$. Gradient calculation by existing PSL methods is summarized in Table 4. After calculating the gradient of $\ell_{\mathrm{psl}}$, the parameter ${\bm{\phi}}$ is iteratively updated as ${\bm{\phi}}\xleftarrow{}{\bm{\phi}}-\eta\frac{\partial\ell_{\mathrm{psl}}}{\partial{\bm{\phi}}}$.
@@ 3.4 Multiobjective Bayesian optimization solvers
PSL-MOBO is an extension of PSL method for expensive MOPs. It optimizes the preference-conditional acquisition function $\alpha({\bm{\theta}}|{\bm{\lambda}})$ over an infinite number of preference vectors to generate a set of promising candidates:
$\min_{\bm{\phi}}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Unif}({\bm{\Delta}}_{m})}\left[\alpha({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}})|{\bm{\lambda}})\right],\qquad{\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}):{\bm{\Delta}}_{m}\xrightarrow{}\mathbb{R}^{N}.$ (7)
Table 5: Supported preference-conditional acquisition functions. $p({\bm{y}}|{\bm{\theta}},\mathcal{D})=\mathcal{N}\left(\hat{{\bm{\mu}}}({\bm{\theta}}),\hat{\bm{\sigma}}^{2}({\bm{\theta}})\right)$ represents the posterior distribution over ${\bm{y}}={\bm{L}}({\bm{\theta}})$ obtained from GP models, $\beta$ is a parameter to balance exploration and exploitation, ${\bm{z}}^{*}=(z_{1}^{*},\ldots,z^{*}_{m})^{T}$ denotes a utopian point of the given MOP, and $\bm{\xi}=(\xi_{1},\ldots,\xi_{m})^{T}$ is a vector induced by the modified Tchebycheff scalarization.
preference-conditional acquisition functions
Tch-LCB $\alpha_{\mathrm{{TchLCB}}}({\bm{\theta}}|{\bm{\lambda}})=\underset{1\leq i\leq m}{\max}\{\lambda_{i}(\hat{\mu}_{i}({\bm{\theta}})-\beta\hat{\sigma}_{i}({\bm{\theta}})-z_{i}^{*})\}$
@@
Currently, our library supports two representative preference-conditional acquisition functions: the Tchebycheff scalarization of lower confidence bound (Tch-LCB) [52, 26, 53] and the expected direction-based hypervolume improvement (DirHV-EI) [51]. We note that DirHV-EI can be regarded as an unbiased estimation of a weighted expected hypervolume improvement. Our library also supports DirHV-EGO, which employs a finite set of predetermined reference vectors as in [33].

### v3

@@ 3.2 MOO solvers
A straightforward way is to use some aggregation functions $g_{\bm{\lambda}}():\mathbb{R}^{m}\mapsto\mathbb{R}$ to convert a MOP to a single objective optimization problem. The reason that optimizing this converted single objective optimization problem will yield Pareto optimal solutions is due to the following two theorems.
Theorem 1 (Adapted from Theorem 2.6.2 [36]).
If $g_{\bm{\lambda}}(\cdot)$ is strictly decreasing w.r.t vector ${\bm{L}}({\bm{\theta}})$, i.e., $g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}^{(a)}))<g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}^{(b)}))$ when ${\bm{L}}_{i}({\bm{\theta}}^{(a)})\preceq_{\mathrm{strict}}{\bm{L}}_{i}({\bm{\theta}}^{(b)})$, then the optimal solution $\theta^{*}$ of $g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))$ serves as a Pareto optimal solution for the original MOP.
Proof. See Mitten’s book [36], Page 22. Similarly, for only decreasing aggregation functions, we have the following theorem.
Theorem 2.
If $g_{\bm{\lambda}}()$ is decreasing w.r.t. vector ${\bm{L}}({\bm{\theta}})$ (i.e., $g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}^{(a)}))\leq g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}^{(b)}))$ when ${\bm{L}}({\bm{\theta}}^{(a)})\preceq_{\mathrm{strict}}{\bm{L}}({\bm{\theta}}^{(b)})$, then the optimal solution $\theta^{*}$ of $g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))$ serves as a weakly Pareto optimal solution for the original MOP.
Proof. Similar to the proof of Mitten’s book [36], Page 22.
In Theorem 2, weakly Pareto optimality means that for a solution ${\bm{\theta}}^{(a)}$, no other solutions ${\bm{\theta}}^{\prime}$ can strictly dominate it, i.e., ${\bm{L}}_{i}({\bm{\theta}}^{\prime})<{\bm{L}}_{i}({\bm{\theta}}^{(a)})$ for all $i\in[m]$. Some common aggregation functions include the linear scalarization (LS) function, where $g_{\bm{\lambda}}^{\mathrm{LS}}({\bm{L}}({\bm{\theta}}))=\sum_{i=1}^{m}\lambda_{i}L_{i}({\bm{\theta}})$, the Tchebycheff function, where $g_{\bm{\lambda}}^{\mathrm{Tche}}({\bm{L}}({\bm{\theta}}))=\max_{i\in[m]}\lambda_{i}\cdot(L_{i}({\bm{\theta}})-z_{i})$ (${\bm{z}}$ is a reference point), Penalty-Based Intersection (PBI) function [29], and COSMOS function [33]. For expressions and other aggregation functions, please refer to Section A.2.
For a preference vector ${\bm{\lambda}}\succ\bm{0}$, $g_{\bm{\lambda}}^{\mathrm{LS}}(\cdot)$ is a strict decreasing function meaning that directly optimizing $g_{\bm{\lambda}}^{\mathrm{LS}}(\cdot)$ yields Pareto optimal solutions (by Theorem 1). However, for any ${\bm{\lambda}}\in{\bm{\Delta}}_{m}$, optimizing $g_{\bm{\lambda}}^{\mathrm{Tche}}(\cdot)$ only yields weakly Pareto optimal solutions (by Theorem 2). For an improper setting of the weight factor $\mu$ (see Section A.2 item 1 and 6), the optimal solution of $g_{\bm{\lambda}}^{\mathrm{PBI}}(\cdot)$ or $g_{\bm{\lambda}}^{\mathrm{COSMOS}}(\cdot)$ can be non-(weakly) Pareto optimal solutions of the original MOP.
An aggregation function is optimized by gradient descent in LibMOON via backpropagation, i.e, ${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\eta{\bm{d}}_{k}={\bm{\theta}}-\eta\frac{\partial g_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}))}{\partial{\bm{\theta}}}|_{{\bm{\theta}}_{k}},$ where ${\bm{d}}_{k}$ is called the updating direction at the $k$th iteration.
Gradient manipulation-based methods.
Besides, directly optimizing the aggregation function, a number of so-called “gradient manipulation methods” solve an updating direction ${\bm{d}}_{k}$ using gradient information for each iteration for some specific purpose. For example, as listed in Table 3, EPO [16] aims to find “exact Pareto solutions” (intersection points of Pareto front and preference vectors), HVGrad [37] aim to maximize the hypervolume of a set of solutions, MOO-SVGD [17] aim to find diverse solutions, and PMTL [13] aim to find sector-constrained Pareto solutions.
Table 3: MOO solvers, properties, and complexities.
Method Solution Property Complexity Pref.
EPO [16] Exact solutions $O(m^{2}nK)$
HVGrad [37] Solutions with maximal HV $O(m^{2}nK^{2})$
MGDA-UB [12] Random solutions $O(m^{2}nK)$
MOO-SVGD [17] Diversity by particles repulsion $O(m^{2}nK^{2})$
PMGDA [38] Solutions under specific demands $O(m^{2}nK)$
PMTL [13] Solutions in sectors $O(m^{2}nK^{2})$
Random [39] Random solutions $O(m^{2}nK)$
Agg-LS [36] Convex part of a PF $O(mnK)$
Agg-Tche [29] Exact solutions $O(mnK)$
Agg-mTche [40] Exact solutions $O(mnK)$
Agg-PBI [29] Approximate exact solutions $O(mnK)$
Agg-COSMOS [33] Approximate exact solutions $O(mnK)$
Agg-SmoothTche [22] Approximate exact solutions $O(mnK)$
  • $m$: number of objectives. $n$: number of decision variables. $K$: number of subproblems. $m$ is usually small (e.g., 2-4), $K$ is relatively large (e.g., 20-40), and $n$ is particularly large (e.g., 10,000). Therefore, $m^{2}$ is not a big concern, while $K^{2}$ and $n^{2}$ are big concerns. Complexity is for time complexity, and Pref. denotes whether this method is preference-based or not.
Interestingly, until now, all these gradient manipulation methods can be implemented in two steps: ① calculating a dynamic weight vector $\tilde{\bm{\alpha}}$ and then ② performing backpropagation on a generalized aggregation function $\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}_{k}))$, where $\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}_{k}))=\sum_{i=1}^{m}\tilde{\alpha}_{i}L_{i}({\bm{\theta}})$. At each iteration, gradient manipulation methods can be equivalently expressed as updating the gradient of its induced generalization aggregation function $\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}_{k}))$. The weight vector $\tilde{\bm{\alpha}}$ are achieved by solving a linear programming (LP) problem (e.g., [16][Eq. 24]) in EPO, a quadratic programming (QP) problem (e.g., [13][Eq. 14]) in PMTL, or other more complex algorithms as used to calculate the hypervolume gradient [41].
Some MOO solvers accept preference vectors ${\bm{\lambda}}$ as input, termed preference-based, affecting Pareto solution positions. The others, called preference-free, do not accept preferences, such as those maximizing dominated hypervolume. A summary of these solvers is in Table 3.
@@
For the previous discussions, we assume that all gradients of objective functions $\nabla L_{i}({\bm{\theta}})$ can be easily achieved via backward propagation. However, for some black-box optimization problems, $\nabla L_{i}({\bm{\theta}})$’s may not easily be achieved. Therefore, LibMOON not only supports first-order optimization, but also supports zero-order optimization methods with estimated gradients ^$\hat{\nabla}L_{i}({\bm{\theta}})$ such as evolutionary strategy (ES) [42].
@@ 3.3 Pareto set learning solvers
Method Property Vector $\tilde{\bm{\alpha}}$ Matrix ${\bm{B}}$
EPO-based PSL [18] Exact solutions MOO solvers BP
PMGDA-based PSL [38] Solutions under specific demands MOO solvers BP
Aggregation-based PSL [18] Solutions with optimal aggregation values BP BP
Evolutionary PSL [43] Mitigating local minima by ES BP ES
LoRA PSL [44, 21, 45] A lighter Pareto model structure BP BP
  • BP: backward propagation, ES: evolutionary strategy.
@@
LibMOON supports Pareto Set Learning (PSL), which trains a model with parameter ${\bm{\phi}}$ to approximate the entire PS/PF. A Pareto model is denoted as ${\bm{\theta}}_{\bm{\phi}}(\cdot):{\bm{\Delta}}_{m}\mapsto\mathbb{R}^{n}$ with input as a preference vector and output as a Pareto solution.
(a) Synthetic problem
(b) MultiTask Learning
Figure 2: Architecture of Pareto models.
PSL Architecture. Pareto models vary in structure. For synthetic problems, the simplest model is a fully-connected neural network that takes a preference as input and outputs the corresponding Pareto solution. In multitask learning, the input is a pair $({\bm{x}},{\bm{y}})$ from dataset $\mathcal{D}$, and the decision variable ${\bm{\theta}}$ represents the target network’s parameter, with $\phi$ as the hypernetwork’s parameter [46]. The loss vector is calculated as ${\bm{L}}({\bm{\theta}})=\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}(\ell(\bm{t}_{\bm{\theta}}({\bm{x}}),{\bm{y}}))$, where $\ell$ is a basic loss function like cross-entropy or mean square error. Structures of these two models are illustrated by Figure 2. Besides these two models, LibMOON also supports LoRA (low rank adaptation)-based PSL [44, 21, 45], which admits a low rank adaptation structure and other structures. PSL structures are decoupled from the training loss and used as plug-ins.
PSL Training. Goal of PSL is to find a model with parameter ${\bm{\phi}}$ optimizing the PSL loss $\ell_{\mathrm{psl}}$ $\min_{\bm{\phi}}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Dir}(\bm{p})}\tilde{g}_{\bm{\lambda}}({\bm{L}}({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}))),$ where $\mathrm{Dir}(\bm{p})$ is a Dirichlet distribution with hyperparameter $\bm{p}$. $\tilde{g}_{\bm{\lambda}}(\cdot)$ can either be a generalized aggregation function as introduced in the previous section or a normal aggregation function. The gradient of $\ell_{\mathrm{psl}}$ can be estimated by the chain rule:
@@
$\underbrace{\frac{\partial\ell_{\mathrm{psl}}}{\partial{\bm{\phi}}}}_{1\times D}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Dir}(\bm{p})}\underbrace{\frac{\partial\tilde{g}_{\bm{\lambda}}}{\partial{\bm{L}}}}_{\tilde{\bm{\alpha}}:(1\times m)}\cdot\underbrace{\frac{\partial{\bm{L}}}{\partial{\bm{\theta}}}}_{{\bm{B}}:(m\times n)}\cdot\underbrace{\frac{\partial{\bm{\theta}}}{\partial{\bm{\phi}}}}_{{\bm{C}}:(n\times D)}.$ (2)
@@
Empirically, the expectation involved in Equation 2 is estimated using a mini-batch of $K$ preferences. The gradient vector $\tilde{\bm{\alpha}}=(\tilde{\alpha}_{1},\ldots,\tilde{\alpha}_{m})$ is computed by MOO solvers as introduced in the previous section. The gradient matrix ${\bm{B}}$ can be calculated either via backpropagation (when gradients are easily obtained) or using a zero-order method such as ES. ${\bm{C}}$ can always be estimated through backward propagation, since ${\bm{\theta}}$ is a continuous vector function of ${\bm{\phi}}$. Gradient calculations in existing PSL methods is summarized in Table 4. Parameter ${\bm{\phi}}$ is iteratively updated by gradient descent: ${\bm{\phi}}\xleftarrow{}{\bm{\phi}}-\eta\frac{\partial\ell_{\mathrm{psl}}}{\partial{\bm{\phi}}}$.
@@ 3.4 Multiobjective Bayesian optimization solvers
PSL-MOBO is an extension of PSL method for expensive MOPs. It optimizes the preference-conditional acquisition function (AF) $\alpha({\bm{\theta}}|{\bm{\lambda}})$ over an infinite number of preference vectors to generate a set of promising candidates: $\min_{\bm{\phi}}\ell_{\mathrm{psl}}=\mathbb{E}_{{\bm{\lambda}}\sim\mathrm{Dir}(\bm{p})}\left[\alpha({\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}})|{\bm{\lambda}})\right],{\bm{\theta}}_{\bm{\phi}}({\bm{\lambda}}):{\bm{\Delta}}_{m}\mapsto\mathbb{R}^{N}.$
Table 5: Supported preference-conditional AFs.
preference-conditional AFs
TLCB $\alpha_{\mathrm{{TLCB}}}({\bm{\theta}}|{\bm{\lambda}})=\underset{i\in[m]}{\max}\{\lambda_{i}(\hat{\mu}_{i}({\bm{\theta}})-\beta\hat{\sigma}_{i}({\bm{\theta}})-z_{i}^{*})\}$
@@
Currently, our library supports two representative preference-conditional AFs: the Tchebycheff scalarization of lower confidence bound (TLCB) [50, 23, 51] and the expected direction-based hypervolume improvement (DirHV-EI) [49]. We note that DirHV-EI can be regarded as an unbiased estimation of a weighted expected hypervolume improvement. Our library also supports DirHV-EGO, which employs a finite set of predetermined reference vectors as in [29].